How much pizza is left if 7 people ate 3/8 each of 4 pizzas?
Answer by stanbon(57219)

How much pizza is left if 7 people ate 3/8 each of 4 pizzas?
--------------------
The 7 people ate 7(3/8) = 21/8 pizza
----------------
The 4 pizzas can be written as 32/8 pizza
----------------
Then 32/8 - 21/8 = 11/8 pizza (amount of pizza that is left)
===================
Cheers,
Stan H.
===================